8. The entry to record a write-off of an uncollectible account when using the direct write off method involves a debit to Allowance for Bad Debts b. debit to Accounts Receivable e. credit to Cash 9. A copy machine costs $45,000 when new and has accumulated depreciation of $44,000. Suppose your company sold the machine for $1,000. What is the result of the disposal transaction a. No gain no loss e. Gain of $1,000 10. Your company sells S 180,000 (selling price) of goods and collects sales of tax of 8%, what current liability does the sales create? d. debits to Bad Debts Expense b. Loss of $1,000 . Loss of $45,000 Sales revenue of $194,400 c. Sales tax payable of $14,400. d.Sales revenue of $180,000 IL (20) The December cash record of ABB Insurance follow b. Unearned Revenue of S194,400 Cash Debit Check No. Cash Credit Dee 3 $ 4,150 2416 5 850 16 2418 2,190 1,870 31 2420 2422 ABB Insurances cash account shows a balance of $16,650 at December 31. On December 31, the company received the following bank statement Bank Statement for December Begianing Balance $14,000 NSF S 700 850 400 160 670 1,960 10 EFT S 400 Dec. 2 4,150 11 Check NO. 2416 19 22 Check NO. 2417 29 Check NO. 2418 31 Check NO. 2419 EFT 16 24 25 560 2,190 1,700 BC SC Ending Balanc$18,590 Nots: BC-bank collection: EFTelferufcient funds check SC-service charge Additional data for the bank reconciliation follows: a. The EFT credit was a receipt of rent. The EFT debit was an insurance payment. b. The NSF check was received from a customer named Dunlap Chemical. c. The $1,700 bank collection was for a note receivable. d. The correct amount of Check 2419 for rent expenise is $1,940 instead of $1,490 on cash records 1) Prepare the bank reconciliation for ABB Insurance as at December 31. 2) Journalize any required entries from the bank reconciliation.
Answer #1

8

Debit to Bad debt Expense

Corrent Entry

pQR8VIHipyWsrbI6WPw5sP8Y8gvyX9ZwlBsId0SB

9

Book value =45000-44000=$1000

Hence No gain No Loss which is sales value-Book value =$1000-1000=0

10

Sales Tax Payable $14400

11

35tGvEo5l9wAAAABJRU5ErkJggg==

Debit Credit
Miscellaneous Expense 10
Cash 10
(Bank Charges debited by bank)
Rent Expense 470
Cash 470
(Error in recording Rent expense)
insurance payment 400
Cash 400
(Payment of Insurance btyEFT)
Account Receivable 700
Cash 700
(NSF Check)
Cash 1700
Note Receivable
(Bank Collection ) 1700
Cash 400
Rent 400
(Rent Collection EFT)